OpenStudy (anonymous):

Next # in series: 7 11 19 35 a. 38 b 50 c. 67 d. 73

OpenStudy (anonymous):

Next is 67

OpenStudy (anonymous):

Logic is 7 + 2^2 = 7+4 =11

OpenStudy (anonymous):

11 + 2^3 = 11 + 8 = 19

OpenStudy (anonymous):

19 + 2^4 = 19 + 16 = 35 and so on

OpenStudy (anonymous):

35 + 2^5 = 35 + 32 = 67